SlideServe
Menu
Préférences
© SlideServe · v2026

Probability

hafwen avatar
9 diapositives · 926 Views · Jul 18, 2014

Probability. Shari carlos Mark shawky Nate stocks. Exercise 2.12 – School Absences. Background:

Probability - slide 1
01 / 9Probability
1 / 17

Télécharger la présentation

Probability
Une image / un lien ci-dessous est fourni (tel quel) pour télécharger la présentation

Politique de téléchargement : le contenu du site vous est fourni TEL QUEL pour votre information et votre usage personnel ; il ne peut être vendu, concédé sous licence ni partagé sur d’autres sites sans l’accord de son auteur.Le contenu vous est fourni TEL QUEL, pour votre information et votre usage personnel uniquement. Téléchargez la présentation en cliquant sur ce lien. Lors du téléchargement, si vous ne parvenez pas à télécharger une présentation, l’éditeur a peut-être supprimé le fichier de son serveur.Pendant le téléchargement, si vous ne parvenez pas à obtenir une présentation, le fichier a peut-être été supprimé par l’éditeur.

F I N

Plus de contenu associé

Transcription de la présentation
  1. Probability Shari carlos Mark shawky Nate stocks

  2. Exercise 2.12 – School Absences • Background: • Data collected at elementary schools in DeKalb County, GA suggest that each year roughly 25% of students miss exactly one day of school, 15% miss 2 days, and 28% miss 3 or more days due to sickness.

  3. Questions • (a) What is the probability that a student chosen at random doesn't miss any days of school due to sickness this year? • (b) What is the probability that a student chosen at random misses no more than one day? • (c) What is the probability that a student chosen at random misses at least one day? • (d) If a parent has two kids at a DeKalb County elementary school, what is the probability that neither kid will miss any school? Note any assumption you must make to answer this question. • (e) If a parent has two kids at a DeKalb County elementary school, what is the probability that both kids will miss some school, i.e. at least one day? Note any assumption you make. • (f) If you made an assumption in part (d) or (e), do you think it was reasonable? If you didn’t make any assumptions, double check your earlier answers.

  4. Answer – a • First, we convert the percentages that we are given into decimals: • 25% is converted to 0.25 • 15% is converted to 0.15 • 28% is converted to 0.28 • We can now calculate the probability that the student didn’t miss any days of school by subtracting the probability that they missed one or more days from 1. • The equation will be set up as follows: • 1 – ((the probability that the student only missed 1 day) + (the probability that the student missed 2 days) + (the probability that the student missed 3 or more days)) • In mathematical terms it would be: • 1 – ((0.25) + (0.15) + (0.28)) = 0.32

  5. Answer - b • To calculate the likelihood that a student missed no more than one day we perform a function similar to Answer A. Except, in this case we want to know whether they missed two or more days. • Set the equation up as follows: • 1 – ((the probability that the student missed 2 days) + (the probability that the student missed 3 or more days) • In mathematical terms: • 1 – ((0.15) + (0.28)) = 0.57

  6. Answer – c • This question asks us the probability that a student misses at least one day. We are given the probability of students missing one day, two days, or three or more. Therefore, in order to answer the question we simply need to add those probabilities together. • Solution: • 0.25 + 0.15 + 0.28 = 0.68

  7. answer – d • This question asks what the probability the neither child misses any school. We apply the multiplication rule for independent processes. • P(A and B) = P(A) x P(B) • In problem A we determined the probability that a random student doesn’t miss any days of school to be 0.32. We apply this to our current question • P(ChildA and ChildB) = P(ChildA) x P(ChildB) • =0.32 x 0.32 • = 0.1024

  8. answer – E • This is a similar problem to D and we apply the same logic. In problem C we determined the probability of a randomly selected student missing at least on day to be 0.68. • Again, we apply the multiplication rule for independent processes to calculate our answer. It should look like this: • P(ChildA and ChildB) = P(ChildA) x P(ChildB) • = 0.68 x 0.68 • = 0.4624

  9. Lets Graph!